(1) check whether the function f given by f(x) = 1/(2x-4)^2 for all x belonging to ]-2,2[ is continuous in the interval ]-2,2[. Is it bounded? Justify your answer.
(2) find the values of p and q so that lim { p sinx + x(1-cosx)}/5x^3 as x tends to infinity is 1/6
1
Expert's answer
2020-03-31T11:02:41-0400

(1) Consider a function "f(x)=\frac{1}{(2x-4)^2}, \ x \in ]-2,2[" . The gived function is continuous on interval "]-2,2[" as a ratio of two continuous functions on "]-2,2[".

The function "f(x)" is unbounded in the interval "]-2,2[" , because there exists a sequence "x_n=2-\frac1n, \ n \in \mathbb{N}", in interval "]-2,2[" , such that "f(x_n)=\frac{n^2}{4} \to \infty" for "n \to \infty".


(2) Let consider a function "g(x)=\frac{p\sin x+x(1-\cos x)}{5x^3}=\frac{p\sin x}{5x^3}+\frac{1-\cos x}{5x^2}." Since "p\sin x" and "1-\cos x" are bounded functions in "\mathbb{R}", then "\lim\limits_{x \to \infty}\frac{p\sin x}{5x^3}=0", "\lim\limits_{x \to \infty} \frac{1-\cos x}{5x^2}=0". Therefore "\lim\limits_{x \to \infty} \frac{p\sin x+x(1-\cos x)}{5x^3}=0" for every values of "p" and "q". Consequently, there is no values of parameters "p" and "q" for which "\lim\limits_{x \to \infty} \frac{p\sin x+x(1-\cos x)}{5x^3}=\frac{1}{6}."